### v6

Abstract
The rapid increase in the number of Computed Tomography (CT) scan examinations has created an urgent need for automated tools, such as organ segmentation, anomaly classification, and report generation, to assist radiologists with their growing workload. Multi-label classification of Three-Dimensional (3D) CT scans is a challenging task due to the volumetric nature of the data and the variety of anomalies to be detected. Existing deep learning methods based on Convolutional Neural Networks (CNNs) struggle to capture long-range dependencies effectively, while Vision Transformers require extensive pre-training, posing challenges for practical use. Additionally, these existing methods do not explicitly model the radiologist’s navigational behavior while scrolling through CT scan slices, which requires both global context understanding and local detail awareness. In this study, we present CT-Scroll, a novel global-local attention model specifically designed to emulate the scrolling behavior of radiologists during the analysis of 3D CT scans. Our approach is evaluated on two public datasets, demonstrating its efficacy through comprehensive experiments and an ablation study that highlights the contribution of each model component. Code: https://github.com/theodpzz/ct-scroll.
keywords:
Multi-label classification, Computed-Tomography, Attention Mechanism.1 Introduction
Computed Tomography (CT) provides detailed imaging of the human body, enabling radiologists to thoroughly examine various anatomical regions, identify abnormalities, and guide patient care from initial diagnosis to follow-up [mazonakis_computed_2016]. However, the growing number of CT scans [broder_increasing_2006] and the associated workload for radiologists have created a pressing need for automated methods to assist in analyzing these volumes [chen_recent_2022]. In medical imaging, and particularly in CT scans, substantial progress has been made in leveraging deep learning techniques to support radiologists in tasks such as segmentation [gu_convformer_2022], image restoration [yuan_deep_2023], classification [draelos_machine-learning-based_2021], and more recently, report generation [hamamci_ct2rep_2024]. As illustrated in Figure 1, multi-label anomaly classification from Three-Dimensional (3D) CT volumes remains a challenging task due to the significant variability in the anomalies that need to be detected.
To process volumetric data and extract meaningful visual features, early approaches predominantly relied on 3D convolutional neural networks (CNNs) to capture spatial dependencies within volumetric data effectively [singh_3d_2020]. Alternatively, some studies adopted conventional 2D architectures by treating a volume as a sequence of slices and subsequently fusing the extracted features [draelos_machine-learning-based_2021]. CNNs excel at capturing local spatial features, and their hierarchical structure facilitates the progressive learning of features, from low-level patterns to high-level semantic representations. More recently, attention mechanisms [vaswani_attention_2023], initially introduced in Natural Language Processing, have demonstrated exceptional performance across diverse text-related tasks [touvron_llama_2023]. This paradigm has been adapted to visual data, including 2D and 3D imaging, by representing images as sequences of 1D tokens derived from flattened 2D or 3D patches. In particular, Vision Transformers (ViTs) [dosovitskiy_image_2021] leverage attention mechanisms to model global context by enabling interactions across different regions of an image. This capability is particularly advantageous for applications requiring a comprehensive understanding of global contexts, making ViTs a promising alternative for complex medical imaging tasks. However, the local receptive fields of CNNs limit their ability to capture global contextual information across large 3D volumes, while ViTs can be computationally expensive when applied to high-dimensional volumetric data and often require large-scale pre-training on extensive datasets to achieve competitive performance [hamamci_generatect_2023]. When radiologists analyze a CT scan, they typically navigate through axial slices to have a global understanding of the volume before focusing on specific anatomical regions of interest [goergen_evidence-based_2013]. If an area appears abnormal, the radiologist often revisits the same slices repeatedly, carefully examining the local context to confirm the diagnosis. Inspired by this diagnostic approach and leveraging the strengths of alternating attention [warner_smarter_2024], originally introduced in NLP, we present a novel alternating global-local attention module, termed the Scrolling Block (SB), illustrated in Figure 2. This module integrates both global and local information through a Sliding Window Attention (SWA) [child_generating_2019, beltagy_longformer_2020] mechanism specifically designed for 3D CT volumes. Our contributions are summarized as follows:
We propose a global-local attention model designed to imitate radiology navigation in 3D CT scans, enhancing multi-label anomaly classification while being achievable with limited computational resources (single GPU, -hour training time).
We demonstrate the robustness and generalizability of our approach through comprehensive cross-dataset evaluation on two public 3D chest CT datasets.
We conduct a comprehensive ablation study analyzing feature reduction and aggregation modules, attention field sizes on model performance and computational efficiency. Our analysis demonstrates how varying the spatial extent of attention mechanisms in our global-local module affects the model’s ability to capture multi-scale anomalies in 3D CT volumes.
2 Related Work
2.1 3D Visual Encoder for Medical Imaging
In the domain of 3D feature extraction, significant efforts have been made across various application areas such as remote sensing, robotic manipulation, and autonomous driving [sarker_comprehensive_2024]. In medical imaging, particularly with 3D CT scans, conventional 3D convolutional neural networks have been widely employed for segmentation [ilesanmi_reviewing_2024] and classification [ho_3d-cnn_2021] tasks. However, CNNs fundamentally lack the ability to model long-range dependencies, limiting their capacity for global context understanding [ma_u-mamba_2024]. More recently, the adaptation of Vision Transformers [chen_vit-v-net_2021] and Swin Transformers [yang_swin3d_2023] to 3D volumes has enabled the interaction of visual tokens corresponding to different patches of the volume via self-attention mechanisms. Positional embeddings are used to preserve spatial information, facilitating better understanding of the volume structure. More recently, CT-ViT [hamamci_generatect_2023] was introduced as a 3D-CT Vision Transformer used to generate 3D CT volumes from free-form medical text prompts. Inspired by ViViT [arnab_vivit_2021], this video-vision transformer learns compact latent representations of 3D volumes by leveraging self-attention and causal attention mechanisms to address the challenges posed by CT scans with varying cranio-caudal coverage. However, Vision Transformers require extensive pre-training on large-scale datasets and a high number of parameters to be effective, which limits their practical applicability. To face this issue, CT-Net [draelos_machine-learning-based_2021] proposes grouping consecutive slices of a CT volume into triplets, which are then passed through a ResNet [he_deep_2015] followed by a small 3D CNN to extract a compact vector representation, subsequently fed into a classification head. This approach showcases robust performance in multi-label classification by effectively capturing fine-grained details, while remaining computationally efficient. However, its reliance on a 3D CNN for feature map reduction limits its ability to model long-range dependencies which can be crucial to capture broader anatomical structures. In this work, we introduce a global-local attention module that enables the modeling of both short-range and long-range dependencies accross slices along the z-axis.
2.2 Global and Local Attention
Global Attention. In both Natural Language Processing (NLP) and computer vision, Transformer-based models leverage global attention [luong_effective_2015], where each embedded token interacts with all other tokens through the self-attention mechanism. This allows for comprehensive contextualization, capturing long-range dependencies and integrating global semantic information into the token representations [devlin_bert_2019].
Local Attention. Despite its effectiveness, global attention suffers from quadratic complexity with respect to sequence length, making it computationally expensive for long sequences in NLP [beltagy_longformer_2020]. To address this, local attention mechanisms such as windowed attention were introduced, restricting each token’s receptive field to a local neighborhood, thereby improving efficiency while preserving essential contextual information. In computer vision, local attention has been successfully adapted in models like Swin Transformer [liu_swin_2021], where image patches interact within localized windows. This hierarchical approach enables efficient processing of high-resolution images and enhances the model’s ability to handle objects with varying scales.
Alternating Attention. Recent advancements in large language models (LLMs) [touvron_llama_2023] have demonstrated the benefits of alternating global and local attention to improve efficiency and contextual modeling. For instance, ModernBERT [warner_smarter_2024] integrates architectural innovations inspired by recent LLMs [team_gemma_2024], alternating between global and local attention layers to balance long-range context aggregation with fine-grained local dependencies.
3 Dataset
CT-RATE dataset. We leverage the publicly available CT-RATE dataset [hamamci_foundation_2024] to train and evaluate our proposed method. This dataset comprises 3D non-contrast chest CT scans, annotated with 18 anomalies extracted from radiology reports using a RadBERT classifier [yan_radbert_2022]. The dataset is partitioned as follows: 17,799 unique patients for the train set, 1,314 unique patients for the validation set and 1,314 unique patients for the test set.
Rad-ChestCT dataset. To assess cross-dataset generalization, we extend our evaluation to the Rad-ChestCT test dataset [draelos_machine-learning-based_2021], which consists of 1,344 3D non-contrast chest CT scans annotated with 83 anomalies extracted using a SARLE labeler from radiology reports. Among these anomalies, we evaluate our method on the 16 anomalies shared with the CT-RATE dataset.
Preprocessing. For all experiments and for both datasets, all CT volumes are preprocessed to ensure uniformity and consistent input characteristics across datasets, enabling robust training and evaluation. Each volume is either center-cropped or padded to achieve a resolution of with an in-plane resolution of mm and mm in the z-axis. Hounsfield Unit values are clipped to the range [, ], before scaling to [, ] and normalization using ImageNet statistics [russakovsky_imagenet_2015].
4 Method
When a radiologist navigates along the longitudinal axis of a CT volume [patel_ct_2024], they scroll through axial slices to detect anomalies. Initially, they perform a global assessment to develop a comprehensive understanding of the volume before revisiting specific slices that may contain abnormalities. Upon identifying a potential anomaly, radiologists frequently scroll back and forth across adjacent slices to incorporate local contextual information, refining their assessment by leveraging both global structure and local details. Model implementation is available here. As illustrated by Figure 2, we propose a method that extracts vector representations from triplets of slices and models their interactions using global and local attention blocks. These attention mechanisms are designed to imitate the scrolling behavior of radiologists, capturing global and local contextual relationships across CT axial slices. The extracted features are then fused to predict the presence of anomalies effectively. The model is trained for multi-label classification using a binary cross-entropy loss function [goodfellow_deep_2016].
4.1 Triplet Slices Embedding
Similar to CT-Net [draelos_machine-learning-based_2021], the slices of the initial volume are grouped in triplets, where each triplet consists of three consecutive slices. This results in a 4D tensor with dimensions (). For each triplet (), a feature map is extracted using a ResNet [he_deep_2015] pre-trained on ImageNet [russakovsky_imagenet_2015], noted , and passed through a Global Average Pooling (GAP) layer to obtain a vector representation for the triplet, noted (), such that:
We employ Global Average Pooling instead of a linear projection or a 3D reducing convolutional layer to significantly reduce the total number of trainable parameters while preserving the local information encoded in the feature maps [li_multi-label_2023].
4.2 Scrolling Block
These vector representations , considered as visual tokens associated with the triplet slices, are then fed into a Scrolling Block (SB), denoted as . A Scrolling Block consists of three Transformer encoders [vaswani_attention_2023]. The first encoder, denoted as , employs global self-attention, enabling each token to aggregate information from the entire volume to have a global understanding of the 3D structure by capturing long-range dependencies along the z-axis. The second and third encoders, denoted as and , use Caudal-Cranial and Cranial-Caudal Sliding Window Attention, respectively, emulating the radiologist’s scrolling behavior along the longitudinal axis to focus on local contextual information. For each triplet slice, the corresponding visual token can only interact with visual tokens associated with slices above it (for Caudal-Cranial modeling, Figure 3.b) or below it (for Cranial-Caudal modeling, Figure 3.c) along the longitudinal axis to capture short-range dependencies, as illustrated by Figure 3. Our method leverages both global attention, capturing long-range dependencies across slices, and local attention, refining contextual representations within localized regions. This design effectively models both short- and long-range interactions along the cranial-caudal axis, mirroring the way radiologists navigate through CT scans for clinical assessment. Each Transformer encoder is followed by a residual connection [he_deep_2015], a normalization layer [ba_layer_2016], and a FeedForward Network leveraging GeGLU, a Gated Linear Unit (GLU)-based activation function that has demonstrated consistent empirical improvements over standard activation functions [shazeer_glu_2020]. This Scrolling Block module generates updated visual tokens with a dimension of 512, denoted as , such that:
Aggregation. The resulting vector representations are aggregated through summation and passed to a classification head, implemented as a lightweight Multilayer Perceptron, denoted as , which predicts a logit vector , such as:
5 Implementation Details
The model was trained for 100k steps with a batch size of , using the AdamW optimizer and a cosine scheduler with a warm-up phase of 20,000 steps and a maximum learning rate of . Training was conducted on a GPU with 48GB of memory.
6 Experimental Results
6.1 Quantitative results
Dataset Method AUROC Accuracy F1 Score W. F1 Score Precision CT-RATE Random Predictions 49.88 0.62 49.89 0.31 27.78 0.51 33.13 0.33 49.85 1.12 ViViT† 79.19 0.28 75.95 0.71 49.91 0.28 54.20 0.35 41.48 0.44 Swin3D† 79.94 0.15 75.95 0.25 50.64 0.25 54.68 0.21 42.07 0.56 CT-Net 79.37 0.27 77.37 0.40 51.39 0.50 56.37 0.32 43.51 0.68 3D CNN† 81.47 0.78 77.80 0.37 52.92 1.08 55.68 0.27 45.46 1.47 CT-Scroll (ours) 81.80 0.22 79.49 0.45 53.97 0.21 58.08 0.28 48.34 1.49 Rad-ChestCT Random Predictions 49.68 0.55 50.40 0.32 35.91 0.41 47.72 0.51 51.51 0.75 ViViT† 67.83 0.38 60.22 1.15 48.59 0.97 60.91 1.20 43.88 0.37 Swin3D† 67.29 0.23 60.67 0.60 47.98 0.41 59.80 0.54 44.03 0.45 CT-Net 67.71 0.83 60.05 1.93 47.53 0.93 60.27 0.92 43.38 1.19 3D CNN† 71.13 0.62 61.08 0.60 49.28 0.93 61.42 0.91 43.38 0.41 CT-Scroll (ours) 71.21 0.37 63.02 0.93 48.55 0.54 59.99 0.58 48.35 0.49
We evaluate the model’s performance using standard metrics: AUROC, F1-Score, precision, and accuracy. For classification, we determine the threshold that maximizes the F1-Score for each of the 18 labels on the validation set, as F1-Score is the harmonic mean of precision and recall [rainio_evaluation_2024]. On the test set, we compute the average of each metric across all labels, as well as the weighted average F1 Score (W. F1 Score) based on label frequencies in the test set. Reported mean and standard deviation metrics were computed over five independent runs with different random seeds to ensure robustness. As shown in Table 4, our method achieves an F1-Score of 53.97 (+ over CT-Net) and an AUROC of 81.80 (+ over CT-Net) on the CT-RATE test set. On the Rad-ChestCT test set, CT-Scroll achieves a +3.29% improvement in AUROC over ViViT, a +3.34% increase over Swin3D and a +3.24% increase compared to CT-Net. On the CT-RATE test set, a paired t-test between our method and CT-Net on all metrics yielded p-values below 0.01, demonstrating the statistical significance of these improvements.
6.2 Ablation study
Impact of the Scrolling Block module. To evaluate the effectiveness of the proposed Scrolling Block, we compare its performance against various traditional modules by replacing the Scrolling Block with these alternatives. Table 2 presents the performance of our models and the contribution of each architectural component. Replacing a small 3D convolutional layer [draelos_machine-learning-based_2021] with a Global Average Pooling layer [li_multi-label_2023] for dimensionality reduction yields a +2.39% improvement in AUROC. Incorporating self-attention via Transformer Encoders enables long-range interactions between visual tokens from triplet slices, improving the F1-Score to 53.32, marking a +1.25% increase over the baseline without self-attention. Integrating local attention via a standard Sliding Window Attention mechanism [beltagy_longformer_2020], after an initial global attention module, leads to a +0.64% improvement in AUROC and a +0.62% increase in F1-score compared to the global-attention-only configuration. Introducing local attention limits the interaction between CT scan slices within the same spatial neighborhood, which could enable the model to learn more fine-grained feature representations, ultimately enhancing anomaly classification performance. Finally, incorporating the Scrolling Block leads to an AUROC of 81.80 (+0.68% increase over global-attention-only configuration) and an F1-score of 53.97 (+1.22% increase over global-attention-only configuration). Inference takes approximately 90 milliseconds, making it suitable for clinical practice applicability.
Method Reduction Interactions AUROC F1 Score # Params GPU Mem. FLOPs ViViT - - 73.19 0.28 49.91 0.28 22 8 0.500 Swin3D - - 79.94 0.15 50.64 0.25 28 14 0.905 3D CNN - - 81.47 0.78 52.92 1.08 0.3 26 0.388 CT-Net 3D Conv. None 79.37 0.27 51.39 0.50 15 15 1.344 - Linear Proj. None 80.42 0.51 52.43 0.71 70 15 1.345 - GAP None 81.21 0.40 52.66 0.41 12 15 1.335 - GAP Tr. Enc. (causal att.) 81.45 0.21 52.98 0.44 16 15 1.337 - GAP Tr. Enc. (global att.) 81.25 0.07 53.32 0.22 16 15 1.337 - GAP Tr. Enc. (global + local) 81.77 0.06 53.65 0.39 16 15 1.337 CT-Scroll GAP Scrolling Block 81.80 0.22 53.97 0.21 16 15 1.337
Window size AUROC Accuracy F1 Score Precision Recall 4 81.54 0.09 78.94 0.37 53.47 0.03 46.99 0.48 65.79 0.97 16 81.80 0.22 79.49 0.85 53.97 0.21 48.34 1.49 65.36 1.21 64 81.46 0.14 79.42 0.72 53.37 0.27 47.75 0.84 64.19 1.44
Impact of the Sliding Window Size. Table 3 presents the model’s performance across various window sizes. CT-Scroll with a window size of 16 outperforms the wider window size of 64 by +1.12% in F1-score.
Qualitative results. Figure 4 illustrates CT axial slices with Grad-CAM activation maps [selvaraju_grad-cam_2019], extracted from the ResNet of the triplet slices embedding module, highlighting CT-Scroll’s ability to identify abnormalities from relevant regions.
7 Conclusion
In this work, we introduce CT-Scroll, a hybrid model for multi-label classification from 3D CT Volumes. Specifically, our approach extracts triplet slices via a 2D CNN to capture fine-grained features, followed by an alternating global-local attention module that models both short- and long-range dependencies along the z-axis. CT-Scroll is evaluated on two public datasets, demonstrating improved multi-label classification performance while maintaining computational efficiency. Additionally, we perform an ablation study analyzing different modules for feature reduction and aggregation. Future work could explore the integration of region-specific information to further enhance classification performance, including the investigation of learnable fusion weights for features extracted from different window sizes. Moreover, taking full advantage of the third dimension by incorporating a 3D CNN module could help to preserve spatial continuity in volumetric data.
We acknowledge [hamamci_foundation_2024] for providing the CT-RATE dataset and [draelos_machine-learning-based_2021] for providing the Rad-ChestCT dataset. This work was performed using HPC resources from GENCI–[IDRIS] [Grant No. 103718]. We thank the support team of Jean Zay for their assistance.
Appendix A Impact of weight inflation
[…]